# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

### v1

@@ 1.3. Main Result and Outline of the proof
where $\varphi(x)$ is a nice enough function. We consider the corresponding mild solution to (1.1), diffusively rescaled:
(1.12) $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx).$
It is easy to see that the mean of $\theta_{n}$ solves the heat equation:
@@
$\mathcal{X}_{n}(t,x):=n^{d/2}(\theta_{n}(t,x)-\bar{\theta}(t,x)).$
Our main result shows that $\mathcal{X}_{n}(t,x)$ has a Gaussian limit, as is expected by the CLT-type scaling.
@@
Let $\alpha>d/2$ and $\gamma\in(0,1/2)$. Assume that $Q$ satisfies (1.2). Then for every $\varphi\in C^{\infty}_{c}(\mathbb{R}^{d})$, $\mathcal{X}_{n}(t,x)$ converges in distribution in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$ to $\mathcal{U}(t,x)$, the solution of the following additive stochastic heat equation:
@@
In the case where $Q$ satisfies Assumption 1.1, the same are true, where the limiting equation is given by (1.14) with $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(z)\,\mathrm{d}z$.
@@
The field $\Psi(t,x)$ is equal to $1$ iff $Q$ is divergence-free, see [10], thus the first part of the theorem covers the case where the correlation function is divergence-free. The point is that the divergence-free condition allows for the control of the SPDE, (1.1), even if the correlation function of the noise is not smooth, or compactly supported. In fact, Assumption 1.1 allows for Kraichnan-type noises, corresponding to $g(\xi)=\langle\xi\rangle^{-(d+\zeta)}$, for $\zeta\in(0,2)$. On the other hand, it seems possible Assumption 1.2 is too strong. Indeed, the results of Section 2.1 can be proved under the assumption that $Q$ is differentiable and Lipschitz, see Remark 2.3. Here, we work under Assumption 1.2 in order to use the results of [10], specifically, (1.15). We believe that (1.15) remains true under weaker assumptions on $Q$, but we do not pursue this here in order not to detract from our main result.
@@
Observe that we take as the initial data $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. When $\varphi$ is nonnegative and $\|\varphi\|_{L^{1}}=1$, we can interpret this choice as starting the diffusion (1.3) with an initial condition $X_{0}\sim\varphi(x)\,\mathrm{d}x$. Then, we can write
@@
Here, $\mathbb{E}_{\varphi}$ is the expectation with respect to the law of the diffusion (1.3), with $\varphi(x)\,\mathrm{d}x$ as the initial distribution and conditional on $V$, and $(\tilde{B}_{t})_{t\geqslant 0}$ denotes a Brownian motion with diffusivity $(\kappa+\nu)I_{d\times d}$, with $\varphi(x)\,\mathrm{d}x$ as an initial distribution.
@@
Let us sketch the basic idea of the proof. We make use of the linearity of (1.1) and the white-in-time correlations of the noise to write (1.4) as a stochastic integral. In particular, (1.4) is a martingale. Therefore, to prove Theorem 1.9, we show that (1.4) is tight in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$, and that its quadratic variation converges to the quadratic variation of the martingale part of the mild solution of (1.14). This, combined with the Skorohod representation theorem, will allow us to conclude.
@@
$\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}.$
This is the content of Proposition 2.6, which is proved in Section 2.1. To prove these estimates, we need to control the expectation of $L^{p}$ norms of $\theta_{t}$, for $p>1$. This control is immediate in the case where the noise is divergence-free, as we can use (1.8). To obtain a similar control under Assumption 1.2, we rely on the correlation functions of (1.7). These functions satisfy a closed-form PDE that has a fundamental solution which, in turn, satisfies appropriate heat kernel bounds (see Propositions 2.1 and 2.2). Using these observations, we can obtain the required control of the $L^{p}$ moments of $\theta_{t}$, see Lemma 2.4.
Having these estimates at hand, we can prove tightness of the laws of $\{\mathcal{X}_{n}\}_{n\geq 1}$ in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$ (see Lemma 2.8 and Proposition 2.9). To calculate the limiting covariance, we make use of the pointwise limiting statistics of $\theta_{n}(t,x)$. In particular, under Assumption 1.2, we make use of the result of [10]. More specifically, for a fixed $\varepsilon>0$, we use the bound [^4] Actually, in [10] this is proved in the case where $\phi$ is a delta function, centered at 0, and for all $t\geqslant t_{0}$, for some $t_{0}>0$. A straightforward adaptation of the arguments in [10] can show (1.15) as well.
(1.15) $\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\big{[}|\theta_{n}(t,x)-q_{t}\ast\varphi(x)\Psi(n^{2}t,nx)|^{2}\big{]}\lesssim_{\varepsilon}n^{-\gamma}$
for some $\gamma>0$, and all $t\geqslant\varepsilon$. Here $q_{t}$ is the standard heat kernel on $\mathbb{R}^{d}$.
@@ 2.1. Correlation Functions and a priori estimates
Throughout this section, we work under Assumption 1.2. We consider (1.1), with initial data $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. It will become apparent that to control moments of the $H^{-\alpha}_{x}$ norm of $\theta_{t}-\bar{\theta}_{t}$, we will need to control $L^{p}_{x}$ norms of $\theta(t,x)$. In particular, we seek to prove a bound of the form
@@
To prove this, we utilize the correlation functions of the model. More specifically, for any $p\in\mathbb{N}$, we define the $p$-th correlation function
@@
(2.3) $\mathcal{C}_{p}(x_{1:p})=\begin{pmatrix}(\kappa+\nu)I_{d}&&Q(x_{1}-x_{2})^{T}&&\dots&&Q(x_{1}-x_{p})^{T}\\ Q(x_{2}-x_{1})&&(\kappa+\nu)I_{d}&&\ldots&&Q(x_{2}-x_{p})^{T}\\ \vdots&&\vdots&&\ddots&&\vdots\\ Q(x_{p}-x_{1})&&Q(x_{p}-x_{2})&&\dots&&(\kappa+\nu)I_{d}\end{pmatrix}$.
@@
(2.4) $\partial_{t}\mathcal{S}_{p}(t,x_{1:p})=\operatorname{Tr}(\nabla^{2}\mathcal{C}_{p}(x_{1:p})\mathcal{S}_{p}(t,x_{1:p})),$
with $\mathcal{S}_{p}(0,x_{1:p})=\varphi^{\otimes p}(x_{1:p})$, where we recall that $\varphi$ is the initial data for (1.1). [^5] Here, we adopt the notation $g^{\otimes p}(x_{1:p}):=g(x_{1})g(x_{2})\dots g(x_{p})$, for a function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
We point out that for a matrix $A$, we denote
@@
(2.5) $\langle\theta_{t},g_{i}\rangle=\langle\varphi,g_{i}\rangle+\int_{0}^{t}\langle\theta_{s},(\kappa+\nu)\Delta g_{i}\rangle\,\mathrm{d}s-\int_{0}^{t}\langle\nabla g_{i},\theta_{s}V(\,\mathrm{d}s)\rangle,$
for all $i=1,\dots,p$ Now, for
@@
where $\partial_{x_{i}^{a}}$ denotes the partial derivative with respect to the $a$-th component of $x_{i}$.
@@
$\langle\mathcal{S}_{p}(t,\cdot),G\rangle=\langle\varphi^{\otimes p},G\rangle-\int_{0}^{t}\langle\mathcal{S}_{p}(s,\cdot),\mathcal{C}_{p}\nabla^{2}G\rangle\,\mathrm{d}s,$
where $G(x_{1:p})=g_{1}(x_{1})g_{2}(x_{2})\dots g_{p}(x_{p})$. Arguing by density (as $\theta_{t}$ is smooth in space) concludes the proof. ∎
@@
is a non-divergence form parabolic PDE with smooth, uniformly elliptic coefficients. Similarly to before, this PDE has a fundamental solution $\tilde{G}_{p}(t,y_{1:p},x_{1:p})$. From [9, Remark 5.12], $\tilde{G}_{p}$ satisfies the bound (2.9). By noticing that $G_{p}(t,y_{1:p},x_{1:p})=\tilde{G}_{p}(t,x_{1:p},y_{1:p})$, we conclude the proof. ∎
@@
This bound is used extensively to prove Lemma 2.4, Lemma 2.5 and Proposition 2.6. As such, it is one of the central ingredients of our arguments. We note that this proposition holds under weaker assumptions. Indeed [9] requires the coefficients to have finite Dini mean oscillation. On the other hand, to make sense of the equation using Kunita’s theory, one convenient assumption would be to take $Q$ to be differentiable and Lipschitz. As such, we expect that we can prove Theorem 1.9 under the latter assumption. As mentioned in Remark 1.10, we do not pursue this here to avoid re-proving statements we need from [10], which are proved there under Assumption 1.2.
With the bound (2.9), one can estimate moments of the $L^{2}$ norm of $\theta_{t}$, as the next lemma shows.
@@
Let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and start the equation (1.1) with initial data $\varphi$. Then for all $r\in\mathbb{N}$, we have
$\sup_{t\in[0,T]}\mathbb{E}\big{[}\|\theta_{t}\|_{L^{2}_{x}}^{2r}\big{]}^{1/2r}\lesssim\|\varphi\|_{L^{2}}.$
@@
$\int_{\mathbb{R}^{dr}}\mathbb{E}\bigg{[}\prod_{i=1}^{r}|\theta_{t}(x_{i})|^{2}\bigg{]}\,\mathrm{d}x_{1:r}.$
@@
$\mathcal{S}_{2r}(t,y_{1:2r})=\int_{\mathbb{R}^{2dr}}G_{2r}(t,y_{1:2r},z_{1:2r})\varphi^{\otimes 2r}(z_{1:2r})\,\mathrm{d}z_{1:2r}.$
@@
$|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\int_{\mathbb{R}^{2dr}}q_{ct}^{\otimes 2r}(y_{1:2r}-z_{1:2r})|\varphi^{\otimes 2r}(z_{1:2r})|\,\mathrm{d}z_{1:2r}.$
@@
With this choice, we have
$\mathbb{E}\bigg{[}\prod_{i=1}^{r}|\theta_{t}(x_{i})|^{2}\bigg{]}=|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\prod_{i=1}^{r}q_{ct}\ast|\varphi|(x_{i})^{2}.$
Integrating over $x_{1:r}$, and using the bound $\|q_{t}\ast\varphi\|_{L^{2}_{x}}\leqslant\|\varphi\|_{L^{2}_{x}}$, concludes the proof. ∎
We are going to need a bound on the correlation function of the rescaled solutions $\theta_{n}$. This is given by the following lemma.
@@
Let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and start equation (1.1) with $n^{-d}\varphi(x/n)$ as the initial data. Recall that $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx)$. We have
$|\mathcal{S}^{n}_{2}(t,x,y)|=|\mathbb{E}[\theta_{n}(t,x)\theta_{n}(t,y)]|\lesssim q_{ct+1}(x)q_{ct+1}(y).$
@@
We can write
@@
Taking $t\rightarrow n^{2}t$, $(x,y)\rightarrow(nx,ny)$, and multiplying both sides by $n^{2d}$ concludes the proof. ∎
Finally, we prove an estimate for the moments of negative Sobolev norms of the solution to (1.1). Similar estimates appear in [23]. The difference here is that we do not assume that the correlation function $Q$ is divergence-free. Recall that $\bar{\theta}$ denotes the solution to (1.13). We have
@@
where $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$ is the initial data of (1.1).
Remark 2.7.
As mentioned above, the working assumption in this section is Assumption 1.2. Nevertheless, a similar bound to (2.10) holds under Assumption 1.1 as well. This can be seen by following the same arguments as in the proof of (2.10) but using (1.8) in place of Lemma 2.4. In fact, more generally, under Assumption 1.1 we have
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big{]}^{1/q}\lesssim_{d,\alpha,\nu,q,T}\|\varphi\|_{p}\big{\|}\widehat{Q}\big{\|}_{p/(2-p)}^{1/2},$
for $p\in(1,2]$, $q\geqslant 1$ and $\alpha\in(\frac{d}{2},\frac{d}{2}+1)$. The proof is similar to the proof below, and since we are not going to use this bound, we skip the details.
Proof of Proposition 2.6.
Recall that $P_{t}=\operatorname{e}^{(\kappa+\nu)t\Delta},\,t\geqslant 0$ is the heat semigroup; by (1.10) and the mild form of the heat equation (1.13), we see that the difference $\theta_{t}-\bar{\theta}_{t}$ is nothing but the stochastic convolution
$Z_{t}:=-\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\,\mathrm{d}B_{k}(r).$
We have
(2.11) $\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $=\bigg{[}\mathbb{E}\Big{\|}\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\,\mathrm{d}B_{k}(r)\Big{\|}_{\dot{H}^{-\alpha}}^{2q}\bigg{]}^{1/q}$
$\lesssim_{q}\bigg{[}\mathbb{E}\Big{(}\sum_{k}\int_{0}^{t}\big{\|}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{\|}_{\dot{H}^{-\alpha}}^{2}\,\mathrm{d}r\Big{)}^{q}\bigg{]}^{1/q},$

### v2

@@ 1.3. Main Result and Outline of the proof
where $\varphi(x)$ is a smooth and compactly supported function, independent of $n$. Let $\theta$ be the corresponding solution to (1.1), which is understood in the Itô form (1.7) and is well-posed by Theorem 1.7 (of course $\theta_{0}$ and so $\theta$ depend on $n$, but, with a small abuse of notation, we will omit this dependence, to make notation lighter). We then consider the diffusive scaling for $\theta$:
(1.12) $\theta^{n}(t,x)=n^{d}\theta(n^{2}t,nx).$
Note that $\theta^{n}$ satisfies (1.7), with initial condition $\theta^{n}(0,x)=\varphi(x)$, but with $V$ replaced by $V^{n}(t,x)=nV(n^{2}t,nx)$. In particular, the mean of $\theta^{n}$ solves the heat equation
@@
(1.14) $\mathcal{X}_{n}(t,x):=n^{d/2}(\theta^{n}(t,x)-\bar{\theta}(t,x)).$
Our main result shows that $\mathcal{X}_{n}(t,x)$ has a Gaussian limit, as expected by the CLT-type scaling:
@@
Let $\alpha>d/2$ and $\gamma\in(0,1/2)$. Assume that $Q$ satisfies Assumption 1.2. Then, for every $\varphi\in C^{\infty}_{c}(\mathbb{R}^{d})$, $\mathcal{X}_{n}(t,x)$, as defined in (1.14), converges in distribution in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$ to $\mathcal{U}(t,x)$, the solution of the following additive stochastic heat equation:
@@
In the case where $Q$ satisfies Assumption 1.1, the same is true, where the limiting equation is given by (1.15) with $V_{\rm eff}^{2}=g(0)\Pi$, where $\Pi$ is the Helmholtz-Leray projection.
@@
The field $\Psi(t,x)$ is equal to $1$ iff $Q$ is divergence-free, see [17]; if $Q$ is also integrable, then a slight modification of the proof in Section 2.3 shows that $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(z)\,\mathrm{d}z$ in this case, thus the first part of the theorem covers the case where the correlation function is divergence-free and integrable. The point is that the divergence-free condition allows for the control of the SPDE, (1.1), even if the correlation function of the noise is not smooth, or compactly supported. In fact, Assumption 1.1 allows for Kraichnan-type noises, precisely we can take $g(\xi)=\langle\xi\rangle^{-(d+\zeta)}$, for $\zeta\in(0,2)$. On the other hand, it seems possible Assumption 1.2 is too strong. Indeed, the results of Section 2.1 can be proved under the assumption that $Q$ is twice continuously differentiable, see Remark 2.3. Here, we work under Assumption 1.2 in order to use the results of [17], specifically, the estimate (1.16) below. We believe that (1.16) remains true under weaker assumptions on $Q$, but we do not pursue this here in order not to detract from our main result.
@@
If one assumes that the correlation function $Q$ is divergence-free, Assumption 1.1 is nearly optimal. Indeed, if one restricts even further and takes $g(\xi)=\langle\xi\rangle^{-(d+\zeta)}$ as in the previous remark, then we cannot prove a statement as in Theorem 1.9 when $\zeta=0$. In this case, the Itô-Stratonovich correction is infinite and (1.1) does not make sense. This is still interesting, however, as in $d=2$ it formally corresponds to a white in time noise that spatially ’looks like’ the $2d$ Gaussian Free Field. With this choice of noise, (1.1) is formally scaling critical, so in analogy to recent works on critical SPDEs [5, 6, 7, 9, 16, 19] we expect that a logarithmic correction to the scaling considered here is needed in order to see non-trivial behavior.
Remark 1.12.
Observe that we take as the initial data $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. When $\varphi$ is nonnegative and $\|\varphi\|_{L^{1}}=1$, we can interpret this choice as starting the diffusion (1.3) with an initial condition $X_{0}\sim n^{-d}\varphi(x/n)\,\mathrm{d}x$. Then, we can write
@@
where $\mathbb{E}_{\varphi}$ is the expectation with respect to the law of the diffusion (1.3), with $\varphi(x)\,\mathrm{d}x$ as the initial distribution and conditional on $V$, and $(\tilde{B}_{t})_{t\geqslant 0}$ denotes a Brownian motion with diffusivity $(\kappa+\nu)I_{d}$, with $\varphi(x)\,\mathrm{d}x$ as an initial distribution.
@@
Let us sketch the basic idea of the proof. We make use of the linearity of (1.7) (more exactly, its rescaled version (2.10)) and the white-in-time correlations of the noise to write (1.14) as a stochastic integral. In particular, (1.14) is a martingale. Therefore, to prove Theorem 1.9, we show that (1.14) is tight in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$, and that its quadratic variation converges to the quadratic variation of the martingale part of the mild solution of (1.15). This, combined with the Skorohod representation theorem, will allow us to conclude.
@@
$\|\theta_{t}^{n}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}.$
This is the content of Proposition 2.7. To prove these estimates, we need to control the moments of $L^{2}$ norms of $\theta_{t}^{n}$, which is done in Section 2.1. This control is immediate in the case where the noise is divergence-free, as we can use (1.8). To obtain a similar control under Assumption 1.2, we rely on the correlation functions of (1.7). These functions satisfy a closed-form PDE that has a fundamental solution which, in turn, satisfies appropriate heat kernel bounds (see Propositions 2.1 and 2.2). Using these observations, we can obtain the required control of the $L^{2}$ moments of $\theta_{t}^{n}$, see Lemma 2.4.
Having these estimates at hand, we can prove tightness of the laws of $\{\mathcal{X}_{n}\}_{n\geq 1}$ in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$, see Lemma 2.9 and Proposition 2.10. To calculate the limiting covariance, we make use of the pointwise limiting statistics of $\theta^{n}(t,x)$. In particular, under Assumption 1.2, we make use of the result of [17]. More specifically, for a fixed $\varepsilon>0$, we use the bound from [17, Theorem 1.1] [^5] Actually, in [17] this is proved in the case where $\varphi$ is a delta function, centered at 0, and for all $t\geqslant t_{0}$, for some $t_{0}>0$. A straightforward adaptation of the arguments in [17] can show (1.16) as well.
(1.16) $\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\big[|\theta^{n}(t,x)-q_{t}\ast\varphi(x)\Psi(n^{2}t,nx)|^{2}\big]\lesssim_{\varepsilon}n^{-\gamma}$
for some $\gamma>0$, and all $t\geqslant\varepsilon$, where $q_{t}$ is the standard heat kernel on $\mathbb{R}^{d}$.
@@ 2.1. Correlation Functions and a priori estimates
Throughout this section, we work under Assumption 1.2. We consider (1.7), with initial data $\theta_{0}\in C_{c}^{\infty}(\mathbb{R}^{d})$. It will become apparent that to control moments of the $H^{-\alpha}_{x}$ norm of $\theta_{t}-\bar{\theta}_{t}$, we will need to control $L^{p}_{x}$ norms of $\theta(t,x)$. In particular, we seek to prove a bound of the form
@@
To prove this, we utilize the correlation functions of the model (1.7). More specifically, for any $p\in\mathbb{N}$, we define the $p$-th correlation function
@@
(2.3) $\mathcal{C}_{p}(x_{1:p})=\begin{pmatrix}(\kappa+\nu)I_{d}&&Q(x_{1}-x_{2})^{\ast}&&\dots&&Q(x_{1}-x_{p})^{\ast}\\ Q(x_{2}-x_{1})&&(\kappa+\nu)I_{d}&&\ldots&&Q(x_{2}-x_{p})^{\ast}\\ \vdots&&\vdots&&\ddots&&\vdots\\ Q(x_{p}-x_{1})&&Q(x_{p}-x_{2})&&\dots&&(\kappa+\nu)I_{d}\end{pmatrix}$.
@@
(2.4) $\partial_{t}\mathcal{S}_{p}(t,x_{1:p})=\operatorname{Tr}(\nabla^{2}(\mathcal{C}_{p}(x_{1:p})\mathcal{S}_{p}(t,x_{1:p}))),$
with $\mathcal{S}_{p}(0,x_{1:p})=\theta_{0}^{\otimes p}(x_{1:p})$, where $\theta_{0}\in C^{\infty}_{c}(\mathbb{R}^{d})$ is the initial data for (1.1).
Here, for a matrix $A$, we denote
@@
(2.5) $\langle\theta_{t},g_{i}\rangle=\langle\theta_{0},g_{i}\rangle+\int_{0}^{t}\langle\theta_{s},(\kappa+\nu)\Delta g_{i}\rangle\,\mathrm{d}s-\int_{0}^{t}\langle\nabla g_{i},\theta_{s}V(\,\mathrm{d}s)\rangle,$
for all $i=1,\dots,p$. Now, for
@@
$\langle\mathcal{S}_{p}(t,\cdot),G\rangle=\langle\theta_{0}^{\otimes p},G\rangle-\int_{0}^{t}\langle\mathcal{S}_{p}(s,\cdot),\mathcal{C}_{p}\nabla^{2}G\rangle\,\mathrm{d}s,$
where $G(x_{1:p})=g_{1}(x_{1})g_{2}(x_{2})\dots g_{p}(x_{p})$. Arguing by density (as $\theta_{t}$ is smooth in space by the last assertion of Theorem 1.7) concludes the proof. ∎
@@
is a non-divergence form parabolic PDE with smooth, uniformly elliptic coefficients. Similarly to before, this PDE has a fundamental solution $\tilde{G}_{p}(t,y_{1:p},x_{1:p})$. From [15, Remark 5.12], $\tilde{G}_{p}$ satisfies the bound (2.9). By noticing that $G_{p}(t,y_{1:p},x_{1:p})=\tilde{G}_{p}(t,x_{1:p},y_{1:p})$, we conclude the proof. ∎
@@
This bound is used to prove Lemma 2.4, Lemma 2.6 and Proposition 2.7. As such, it is one of the central ingredients of our arguments. We note that this proposition holds under weaker assumptions. Indeed [15] requires the coefficients to have finite Dini mean oscillation. On the other hand, to make sense of the equation using Kunita’s theory, one convenient assumption would be to take $Q$ to be twice continuously differentiable. As such, we expect that we can prove Theorem 1.9 under the latter assumption. As mentioned in Remark 1.10, we do not pursue this here to avoid re-proving statements we need from [17], which are proved there under Assumption 1.2.
With the bound (2.9), one can estimate moments of the $L^{2}$ norm of $\theta_{t}$. For our purposes, we will need to estimate moments of the rescaled solutions $\theta^{n}$. In particular, we have the following lemma
@@
Let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and start equation (1.7) with $\theta_{0}(x)=n^{-d}\varphi(x/n)$ as the initial data. Recall that $\theta^{n}(t,x)=n^{d}\theta(n^{2}t,nx)$. Then for all $r\in\mathbb{N}$, we have
$\sup_{t\in[0,T]}\mathbb{E}\big[\|\theta^{n}(t)\|_{L^{2}_{x}}^{2r}\big]^{1/2r}\lesssim\|\theta_{0}\|_{L^{2}}.$
@@
$\int_{\mathbb{R}^{dr}}\mathbb{E}\bigg[\prod_{i=1}^{r}|\theta_{t}^{n}(x_{i})|^{2}\bigg]\,\mathrm{d}x_{1:r}.$
@@
$\mathcal{S}_{2r}(t,y_{1:2r})=\int_{\mathbb{R}^{2dr}}G_{2r}(t,y_{1:2r},z_{1:2r})\theta_{0}^{\otimes 2r}(z_{1:2r})\,\mathrm{d}z_{1:2r}.$
@@
$|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\int_{\mathbb{R}^{2dr}}q_{ct}^{\otimes 2r}(y_{1:2r}-z_{1:2r})|\theta_{0}^{\otimes 2r}(z_{1:2r})|\,\mathrm{d}z_{1:2r}.$
@@
With this choice, and recalling that $\theta_{0}(z)=n^{-d}\varphi(z/n)$ we have
$\mathbb{E}\bigg[\prod_{i=1}^{r}|\theta_{t}(x_{i})|^{2}\bigg]$ $=|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\prod_{i=1}^{r}q_{ct}\ast|\theta_{0}|(x_{i})^{2}$
$=\prod_{i=1}^{r}\bigg(\int_{\mathbb{R}^{d}}q_{ct}(x_{i}-nz)|\varphi(z)|\,\mathrm{d}z\bigg)^{2}.$
Taking $t\rightarrow n^{2}t$, $x_{i}\rightarrow nx_{i}$, multiplying both sides by $n^{2dr}$, and integrating over $x_{1:r}$, yields the bound
$\mathbb{E}\big[\|\theta^{n}(t)\|_{L^{2}_{x}}^{2r}\big]=\int_{\mathbb{R}^{dr}}\mathbb{E}\bigg[\prod_{i=1}^{r}|\theta_{t}^{n}(x_{i})|^{2}\bigg]\,\mathrm{d}x_{1:r}\lesssim\|\varphi\|_{L^{2}},$
where we also used the scaling properties of the standard heat kernel and the bound $\|q_{t}\ast\varphi\|_{L^{2}_{x}}\leqslant\|\varphi\|_{L^{2}_{x}}$. This concludes the proof. ∎
Remark 2.5.
More generally, the same argument works to bound moments of the $L^{2}_{x}$ norm of a solution to (1.7) without rescaling. In particular, if $\theta_{t}$ solves (1.7) with $\varphi\in L^{2}(\mathbb{R}^{d})$ as the initial data, then
$\sup_{t\in[0,T]}\mathbb{E}\big[\|\theta(t)\|_{L^{2}_{x}}^{2r}\big]^{1/2r}\lesssim\|\theta_{0}\|_{L^{2}}$
One could use a similar argument to derive bounds on $\mathbb{E}[\|\theta_{t}\|_{L^{p}_{x}}^{pr}]$, for $p,r\in\mathbb{N}$, but we do not pursue this here.
We are also going to need a pointwise bound on the correlation function. This is given by the following lemma.
@@
Under the same setting as in the previous lemma, we have
$|\mathcal{S}^{n}_{2}(t,x,y)|=|\mathbb{E}[\theta^{n}(t,x)\theta^{n}(t,y)]|\lesssim q_{ct+1}(x)q_{ct+1}(y).$
@@
The proof is very similar to the proof of the previous lemma. We can write
@@
Again, taking $t\rightarrow n^{2}t$, $(x,y)\rightarrow(nx,ny)$, and multiplying both sides by $n^{2d}$ concludes the proof. ∎
2.2. Proof of Theorem 1.9: The compressible case
Here, we prove Theorem 1.9, under the Assumption 1.2. As mentioned in the introduction, the idea is the following: First, we will prove that $(\mathcal{X}_{n})_{n}$, defined in (1.14), is tight and then we will characterize the law of all limiting points.
Recall that $\theta$ solves equation (1.7) with initial condition $\theta(0,x)=n^{-d}\varphi(x/n)$; $\theta^{n}$, defined by $\theta^{n}(t,x)=n^{d}\theta(n^{2}t,nx)$, satisfies the equation
(2.10) $\partial_{t}\theta^{n}+\nabla\cdot(\theta^{n}V^{n})=(\kappa+\nu)\Delta\theta^{n},$
with the rescaled initial condition
$\theta^{n}(0,x)=\varphi(x),$
which is independent of $n$. Here $V^{n}$ in (2.10) is defined by $V^{n}(t,x)=nV(n^{2}t,nx)$ and has spatial covariance function $Q^{n}(x)=Q(nx),\,x\in\mathbb{R}^{d}$; therefore, it has the Fourier transform
(2.11) $\widehat{Q^{n}}(\xi)=n^{-d}\widehat{Q}(n^{-1}\xi).$
Note that $V^{n}=\dot{W}^{n}$ and thus by (1.6),
as a result,
by the scaling property of Brownian motions. Finally, note that $Q^{n}(0)=Q(0)=2\nu I_{d}$ is independent of $n$.
To prove Theorem 1.9, we first prove an estimate for the moments of negative Sobolev norms of the solution to (2.10). Similar estimates appear in [33], see also [21] for the estimates in the case of torus. The difference here is that we do not assume that the correlation function $Q$ is divergence-free. Recall that $\bar{\theta}_{t}(x)$ solves (1.13) with $\varphi$ as the initial data.
@@
(2.13) $\mathbb{E}\Big[\sup_{t\in[0,T]}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big]^{1/q}\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big\|\widehat{Q}\big\|_{\infty}^{1/2}n^{-d/2},$
In particular,
$\mathbb{E}\Big[\sup_{t\in[0,T]}\|\mathcal{X}_{n}(t)\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big]^{1/q}\lesssim_{d,\alpha,q}\|\varphi\|_{2}\big\|\widehat{Q}\big\|_{\infty}^{1/2}.$
Proof.
Recall that $P_{t}=\operatorname{e}^{(\kappa+\nu)t\Delta},\,t\geqslant 0$ is the heat semigroup; using the mild form of (2.10) and of the heat equation (1.13), we see that the difference $\theta_{t}^{n}-\bar{\theta}_{t}$ is nothing but the stochastic convolution
$Z_{t}^{n}:=-\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}^{n}\sigma_{k}^{n})\,\mathrm{d}B_{k}^{n}(r),$
where $\sigma^{n}_{k}(x)=\sigma_{k}(nx)$, $B^{n}_{k}(r)=n^{-1}B_{k}(n^{2}r)$, $k\geq 1$ We have
(2.14) $\big[\mathbb{E}\|Z_{t}^{n}\|_{\dot{H}^{-\alpha}}^{2q}\big]^{1/q}$ $=\bigg[\mathbb{E}\Big\|\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}^{n}\sigma_{k}^{n})\,\mathrm{d}B_{k}^{n}(r)\Big\|_{\dot{H}^{-\alpha}}^{2q}\bigg]^{1/q}$
$\lesssim_{q}\bigg[\mathbb{E}\Big(\sum_{k}\int_{0}^{t}\big\|P_{t-r}\nabla\cdot(\theta_{r}^{n}\sigma_{k}^{n})\big\|_{\dot{H}^{-\alpha}}^{2}\,\mathrm{d}r\Big)^{q}\bigg]^{1/q},$